OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $7,000, with an annual interest rate of 9 percent and a term of 1 year. Round to the nearest cent as needed.

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

I AM REALLY SORRY THAT I CANT HELP WITH THIS ONE! WILL THIS SITE HELP? http://www.csgnetwork.com/simpleloansintcalc.html

OpenStudy (anonymous):

darn, I will see. thanks anyways

OpenStudy (anonymous):

LET ME KNOW

OpenStudy (espex):

Well let's see.. Simple add-on interest means that we add the interest so a principle of 7000 * 0.09 = 630. Since the term is only one year we should have a total repayment of 7630. There are 12 payments in the term so you divide your loan by the number of payments, 7630/12= 635.83 a month.

OpenStudy (anonymous):

aww perfect! that's what I got too! would you mind checking my next problem as well?

OpenStudy (anonymous):

Emily is buying a new refrigerator for $1,400. The distributor is charging her an annual interest rate of 10.7% and is using the add-on method to compute monthly payments. If she pays off the loan in 36 months, what are her monthly payments? If she makes a down payment of $220, how much will her monthly payments be? Do not round until the final answer. Then, round to the nearest cent.

OpenStudy (anonymous):

I got $51 and $43

OpenStudy (espex):

For the first part I have a payment of 43.05 and in the second part I get 36.29

OpenStudy (anonymous):

hmm okay, I'll double check it. thanks! wanna try another one out for me? :)

OpenStudy (espex):

I can give it a look.

OpenStudy (anonymous):

A credit card calculates interest using the average daily balance method. The card charges 18.2% annual interest rate on the average daily balance. The following transactions occurred during the June 1 – June 30 billing period.

OpenStudy (anonymous):

Find the average daily balance for the billing period. Round to the nearest cent.

OpenStudy (espex):

Well from 6/1 - 6/5 you have a balance of 2645.38 = 13266.9 from 6/6 - 6/7 you have 1345.38 = 2690.76 6/8 you have 1379.65 6/9 - 6/16 is 1517.82 = 15178.02 6/17 - 6/26 is 1684.70 = 16847.00 6/27 - 6/30 is 1898.75 = 7595.00 For a total of 56957.33/30 = 1898.58 Assuming I did my math correctly, I expect that you will verify all results.

OpenStudy (anonymous):

THANKS